# Incorporating Magnetic Field Characteristics into EUV-Based Automated Segmentation of Coronal Holes

### v1

@@ 3.3 Effects of Magnetic Unipolarity on Segmentation
Following the pipeline established by Grajeda et al. (2023) for the creation of full-scale ($4096\times 4096$ pixels) CH segmentations, we begin by verifying that the header information of the Level 1 193 Å EUV image is correct (and updating it, if needed, using aiapy.calibrate.update_pointing), then converting to level 1.5 images using aiapy.calibrate.register (Barnes et al., 2020). From there, we correct for limb brightening using the method of Verbeeck et al. (2014) and define the initial contour using the same threshold parameter ($\alpha=0.3$). Magnetic field data are prepared by taking the corresponding HMI magnetogram and aligning it to the EUV observation using the reprojection tool, reproject.reproject_interp, of Robitaille et al. (2020). The EUV and HMI data are combined into a two-channel image, with 193 Å EUV data in the first channel ($i=1$) and HMI data in the last ($i=2$). We note here that Ko et al. (2014) calculated unipolarity using an estimate of the radial magnetic field. The results herein presented are for the line-of-sight magnetic field, due to improved stability.
Matching Grajeda et al. (2023), off-disk areas in both channels are set to the mean intensity of the non-CH region. This operation is performed on a per-channel basis, and reset every 10 iterations. Evolution is performed using the same parameters as Grajeda et al. (2023) for contour length ($\mu=0$), and for 193 Å homogeneity ($\lambda^{\rm{H}+}_{1}=1$ and $\lambda^{\rm{H}-}_{1}=1/50$). Additionally, for the EUV data, we set the unipolarity parameters $\lambda^{\rm{U}+}_{1}=\lambda^{\rm{U}-}_{1}=0$. For the HMI data, we set the homogeneity parameters $\lambda^{\rm{H}+}_{2}=\lambda^{\rm{H}-}_{2}=0$, and define the unipolarity parameters as $\lambda^{\rm{U}+}_{2}=1$, and $\lambda^{\rm{U}-}_{2}=1/50$.
The IOU, SSIM, GCE, and LCE between the full-scale 193 Å segmentations of Grajeda et al. (2023) and our EUV+HMI segmentations, organized by CR, are presented in Figure 1. We note that, across all four metrics, CR 2133 displays the largest median discrepancy. Given that filaments are most prevalent in CR 2133, this is consistent with the expectation that the magnetic unipolarity parameters are reducing filament contamination. Visual inspection of the segmentation pairs with the lowest IOU for CR2133, of which four samples are provided in Figure 2, confirms this speculation by showing that filament regions, while still present, are consistently diminished to a small region surrounding the initial seed.
@@
Figure 1: Comparison between full-scale segmentation results using only 193 Å data, and using both 193 Å and HMI magnetogram data, organized by CR. The box outlines the range between Q1 and Q3, with the median value in orange. The whiskers show 1.5 times the interquartile range. Outliers are marked with circles.
@@
(d) IOU:0.30823
Figure 2: Example of full-scale($4096\times 4096$ pixels) segmentation results using only 193 Å observations (center), and using both 193 Å and HMI magnetogram observations (right) over CR 2133. In each figure the leftmost image is the 193 Å observation. The title of each figure is the record time for the EUV data.
Referring back to Figure 1, we note that that CR 2099 shows a large range between Q1 and Q3 in IOU, GCE, and LCE, but not in SSIM. We also note that CR 2099 shows the second largest median discrepancy in IOU, but not in the remaining three metrics. Both of these results are consistent with a change in overall CH area while retaining the overall form or shape of the identified CHs. Grajeda et al. (2023) note that CR 2099 had the largest number of segmentations that changed from targeting CHs to QS as a result of including excessive QS in $C^{+}$. This may suggest that our new formulation is reducing the growth of $C^{+}$ into QS regions, which is another motivation for inclusion of magnetic field into evolution rather than just post-hoc processing. Visual inspection of the segmentation pairs with the lowest IOU for CR 2099, of which four samples are provided in Figure 3, also confirms this speculation.
@@
(d) IOU:0.34602
Figure 3: Example of full-scale ($4096\times 4096$ pixels) segmentation results using only 193 Å observations (center), and using both 193 Å and HMI magnetogram observations (right) over CR 2099. In each figure the leftmost image is the 193 Å observation. The title of each figure is the record time for the EUV data.
In summary, incorporation of magnetic field information into ACWE (via a unipolarity metric operating on the HMI data) appears to have had negligible impact on the segmentation of CHs. At the same time, incorporation of the magnetic field appears to have minimized area of detected filaments and alleviated many cases of change of target. We will return to both of these results in Section 4 after first considering effects of spatial resolution on segmentation accuracy.
@@ 3.4 Effects of Spatial Resolution on Segmentation Accuracy
As noted in Grajeda et al. (2023), the iterative nature of ACWE results in segmentations at full-scale ($4096\times 4096$ pixels) taking several minutes to complete. Reducing the image to one-eighth-scale resolution ($512\times 512$ pixels) reduces segmentation time to a matter of seconds. In EUV-only CH segmentations, Grajeda et al. (2023) note that this process results in highly similar segmentations.
In order to determine the viability of using one-eighth-scale resolution images for EUV+HMI magnetogram segmentations, we replicate the pipeline of Grajeda et al. (2023) for producing segmentations at this reduced spatial resolution. This is done by reducing the scale of the 193 Å image and the corresponding HMI magnetogram to $512\times 512$ pixels using the scikit-image function skimage.transform.resize (van der Walt et al., 2014) after aligning the observations, but before correcting for limb brightening in EUV and before defining the seed using the EUV observation. We then evolve the initial seed on the decimated images, using the same weights as Section 3.3. After each segmentation has converged, it is upscaled back to the full $4096\times 4096$ pixel resolution and compared with the EUV+HMI magnetogram segmentations developed in Section 3.3.
@@
Figure 4: Comparison between full-scale ($4096\times 4096$ pixels) segmentation results using both 193 Å and HMI magnetogram data, and one-eighth-scale ($512\times 512$ pixels) segmentation results using both 193 Å and HMI magnetogram data for CR 2133. The box outlines the range between Q1 and Q3, with the median value in orange. The whiskers show 1.5 times the interquartile range. Outliers are marked with circles.
The effects of reducing spatial resolution on segmentation similarity for CR 2133 are summarized in Figure 4. Likewise, the effects of reducing spatial resolution on segmentation similarity for CRs 2099-2101 are summarized in Figure 5. In both figures, we provide similarity for bi-linear interpolation, noting that five additional interpolation methods (nearest-neighbor, bi-quadratic, bi-cubic, bi-quartic, and bi-quintic interpolation) were also tested and yielded similar results. We note here that the results in Figure 5 are consistent with the results in Grajeda et al. (2023), and that the results in Figure 4 are improved, suggesting that decimation to one-eighth-scale resolution remains a viable method for minimizing algorithm runtime despite (and in some cases, because of) the addition of decimated magnetogram data.
@@
Figure 5: Comparison between full-scale ($4096\times 4096$ pixels) segmentation results using both 193 Å and HMI magnetogram data, and one-eighth-scale ($512\times 512$ pixels) segmentation results using both 193 Å and HMI magnetogram data for CRs 2099-2101. The box outlines the range between Q1 and Q3, with the median value in orange. The whiskers show 1.5 times the interquartile range. Outliers are marked with circles.
Visual inspection revealed the same three discrepancies between full- and reduced-resolution segmentations as in Grajeda et al. (2023): absence of smaller regions due to the downsampling process removing the dark pixels that formed the initial seed, the presence or absence of spurious bright regions at different scales, and reduced fidelity along the contour boundary due to the lower spatial resolution. All three effects are demonstrated in Figure 6. No new discrepancies appear to have been added by introducing spatially downsampled HMI magnetogram data.
Figure 6: Example of full-scale ($4096\times 4096$ pixels) and one-eighth-scale ($512\times 512$ pixels) segmentation both generated using 193 Å and HMI magnetogram data. The title of the figure is the record time of the EUV data.
The first of the three effects, absence of smaller regions due to the downsampling process removing the dark pixels that formed the initial seed, is especially prevalent in the case of filaments, which, due to their thin structure, are more likely to lose dark pixels that form the initial seed when spatially downsampled. This effect explains the larger discrepancy that Grajeda et al. (2023) found in full- and reduced-scale segmentations for CR 2133. In particular, the fact that addition of HMI magnetogram data constrained filament evolution at the full-scale segmentation ensured higher similarity between the filament-absent one-eighth-scale segmentations, and the filament-present full-scale segmentations.
(a) Intersection Over Union
(b) Structural Similarity
(c) Global Consistency Error
(d) Local Consistency Error
@@
Figure 7: Comparison between full-scale ($4096\times 4096$ pixels) segmentation results using both 193 Å and HMI magnetogram data, and one-eight-scale ($512\times 512$ pixels) segmentation results using both 193 Å and HMI magnetogram data for CR 2133 when the full-Scale seed is utilized to generate the one-eighth-scale segmentation. The box outlines the range between Q1 and Q3, with the median value in orange. The whiskers show 1.5 times the interquartile range. Outliers are marked with circles.
@@
Figure 8: Effects of seed transfer on EUV+HMI magnetogram segmentations generated at one-eighth-scale ($512\times 512$ pixels) resolution (right), compared to the corresponding full-scale segmentations (center). In each figure the leftmost image is the 193 Å observations. The title of each figure is the record time for the EUV data.
@@
Due to the aforementioned absence of some filaments in CR 2133 at one-eighth-scale resolution, we additionally studied the effects of transferring the seed (initial contour) generated at full-scale resolution to the one-eighth scale resolution. This is done to ensure that all filaments present in the full-scale segmentations are also present in the one-eighth scale segmentations, and therefore to verify that the spatially-decimated HMI magnetogram data still proves effective at correctly constraining the CH segmentations. Due to differing spatial resolutions ($4096\times 4096$ pixels at full scale and $512\times 512$ pixels at one-eighth scale), the full-scale seed must be downsampled to the new resolution. Since every pixel at the one-eighth-scale resolution represents an $8\times 8$ pixel patch of the original image, in order to ensure that even a region represented with only one pixel in the full-scale seed is retained at the one-eighth-scale resolution, we apply dilation using a $4\times 4$ pixel square using the scikit-image function skimage.morphology.dilation (van der Walt et al., 2014) prior to downsampling the initial seed. The effects of this process are summarized in Figure 7. We note that this process had minimal effect on SSIM, GCE, and LCE, but further improved IOU, suggesting that the added filaments were constrained in the same manner as before. This is confirmed through a visual inspection of the data. We provide, for reference, Figure 8, which provides the full-scale and one-eighth-scale segmentations generated utilizing the seed-transfer process for the same four examples from Figure 2. Transferring the seed for the remaining three CRs had no appreciable effect.
@@
In summary, reducing the resolution of HMI magnetograms via standard image resizing procedures does not appear to affect the segmentation performance. We note that this image decimation does not take into account magnetic flux conservation (i.e., correction for differing areas of pixels across the disk), but appears to retain statistics relevant to measures of unipolarity. Additionally, we find that a loss of seeds (individual regions in the initial contour) related to filaments does not appear to be the largest contributing source of this performance. Rather, it appears that reduced-resolution HMI data retains relevant information for constraining evolution of the contour, reducing the contribution of filaments. Since filaments are still included in the segmentations, albeit with reduced area, we now consider incorporation of magnetic field information into the seeding process to further reduce filament contamination. This reinforces the importance of incorporating magnetic field information throughout the process of CH segmentation rather than as a post-hoc process.
@@ 4 Incorporation of Magnetic Field Data into the Seeding Process
Due to improved runtime and minimal effect on segmentation quality, we use the one-eighth-scale segmentations for the second step (iterative refinement of the contour) in ACWE, even when utilizing HMI magnetogram data. Additionally, due to reduced filament contamination in the initial contours generated at one-eighth-scale resolution, we seed the algorithm after spatially downsampling the EUV and HMI magnetogram observations. With that noted, we recognize that while this formulation has improved the quality of the segmentations generated using ACWE, filament contamination, though minimized, persists. This section, therefore, explores the use of magnetic field data in initial seeding in order to further reduce filament contamination.
@@ 4.1 Complications with Estimating Unipolarity from Initial Seeds
Predicting the unipolarity of a final evolved region through direct evaluation of the seed, regardless of the scale at which seeding occurred, is not possible due to two issues. First, multiple seeds (individual regions in the seed/initial contour) often contribute to a single CH. Second, these seeds are often small regions, and are therefore vulnerable to the effects of small-scale statistics which results in erroneous estimations of unipolarity at the seeding stage. In order to demonstrate this, we will briefly return to the EUV-only segmentations for the following two studies.
@@
First, regardless of scale, multiple regions within the initial seed often contribute to a single CH or filament detection. The effects of this are partially demonstrated in Figure 2, where the use of HMI magnetogram data minimized evolution of filament regions, resulting in small areas around the initial seeds rather than the unified region seen in the EUV-only segmentations of Grajeda et al. (2023). Further demonstration of this behavior, however, can be shown by counting the number of regions present in the initial seed and comparing that to a count of the number of regions present in the final EUV-only segmentation. This comparison for CR 2133 is provided in Figure 9, for the full-scale EUV-only segmentations, and Figure 10, for the one-eighth-scale EUV-only segmentations. In both figures the dashed-blue line represents an ideal case where for every region in the initial seed, one region (CH or filament) is produced in the final segmentation. Any cases above this line would represent a case where there were more regions in the final segmentation (region splitting), while any case below this represents an example of more regions in the initial seed compared to the final contour (region merging). We note that region splitting does not occur, however region merging is prevalent within the data.
@@
Figure 9: Scatter plot showing the number of regions in the initial seed (x-axis) and number of regions in the final segmentation (y-axis) for EUV-only segmentation at full scale ($4096\times 4096$ pixels) for CR 2133. The color of each point denotes the number of images in the dataset that had that specific number of input regions versus that specific number of output regions. The diagonal line is the $x=y$ line, representing the case where for every region in the initial seed there is exactly one region in the final segmentation.
@@
Figure 10: Scatter plot showing the number of regions in the initial seed (x-axis) and number of regions in the final segmentation (y-axis) for EUV-only segmentation at one-eighth scale ($512\times 512$ pixels) for CR 2133. The color of each point denotes the number of images in the dataset that had that specific number of input regions versus that specific number of output regions. The diagonal line is the $x=y$ line, representing the case where for every region in the initial seed there is exactly one region in the final segmentation.
@@
Multiple seeds contributing to a single CH is only problematic due to the fact that it exacerbates the second issue: small-scale statistics. It is reasonable to assume that regions in the initial contour with close physical proximity are part of the same final region, however, we note that regions with a large physical distance between them can belong to the same solar feature. In the counts for Figure 9 and Figure 10, we account for proximity by defining a region in the same manner as we did in Grajeda et al. (2023), namely a region, whether in the initial seed or in the final segmentation, consists of all pixels that will become a single eight-connected component when dilated with either a $40\times 40$ pixel square at full scale, or a $5\times 5$ pixel square at one-eighth scale. This same definition is applied for the plots in Figure 11 (full scale) and Figure 12 (one-eighth scale) which show scatter plots of initial versus final unipolarity for each region in the EUV-only segmentations.
@@
Figure 11: Scatter plot showing the line-of-sight unipolarity of the initial seeds (x-axis) versus the line-of-sight unipolarity of corresponding region in the final segmentation (y-axis) for EUV-only segmentation at full scale ($4096\times 4096$ pixels) for all seed regions in CR 2133.
@@
Figure 12: Scatter plot showing the line-of-sight unipolarity of the initial seeds (x-axis) versus the line-of-sight unipolarity of corresponding region in the final segmentation (y-axis) for EUV-only segmentation at one-eighth-scale ($512\times 512$ pixels) for all seed regions in CR 2133.
@@
We note that both Figure 11 and Figure 12 indicate regions with a high final (y-axis) value across all initial (x-axis) unipolarity values. This indicates that bipolar regions may appear to have any unipolarity, including high unipolarity, when evaluating the initial seed. We also note, referring to the cluster along the y-axis of Figure 11, that regions reporting high unipolarity at the initial seed are especially prevalent at the full-scale resolution. This explains why the initial seeds in filament regions remain, even when evolving using both EUV and HMI magnetogram observations. The significantly smaller number of observations on the y-axis of Figure 12 provides further indication that seeding at the reduced one-eighth-scale resolution helps to minimize filament contamination.
@@
In summary, we have demonstrated that multiple seeds contribute to a final evolved region (be it a CH or a filament). Furthermore, we have shown that initial seeds may provide an inaccurate estimate of the unipolarity of the final region, mainly due to small-sample statistics. In light of these results, we now consider a means to remove regions associated with filaments from the ACWE evolution.

### v2

@@ 3.3 Effects of Magnetic Unipolarity on Segmentation
Following the pipeline established by Grajeda et al. (2023) for the creation of full-scale ($4096\times 4096$ pixels) CH segmentations, we first verify that the header information of the Level 1 193 Å EUV image is correct (and update it, if needed, using aiapy.calibrate.update_pointing). Next we convert to Level 1.5 images using aiapy.calibrate.register (Barnes et al., 2020). Finally, we correct for limb brightening using the method of Verbeeck et al. (2014) and define the initial contour using threshold parameter $\alpha=0.3$. Magnetic field data are prepared by aligning the corresponding HMI magnetogram to the EUV observation using reproject.reproject_interp (Robitaille et al., 2020). EUV and HMI data are combined into a two-channel image, with 193 Å EUV data in the first channel ($i=1$) and HMI data in the last ($i=2$). We note that Ko et al. (2014) calculated unipolarity using an estimate of the radial magnetic field, but we use the LOS field, due to improved stability.
As in Grajeda et al. (2023), off-disk areas in both channels are set to the mean intensity of non-CH regions on a per-channel basis and reset every 10 iterations. QUACK evolution is performed using the same parameters as Grajeda et al. (2023) for contour length ($\mu=0$), and for 193 Å homogeneity ($\lambda^{\rm{H}+}_{1}=1$, $\lambda^{\rm{H}-}_{1}=1/50$). We set unipolarity parameters $\lambda^{\rm{U}+}_{1}=\lambda^{\rm{U}-}_{1}=0$ for the EUV data. For the HMI data, we set homogeneity parameters $\lambda^{\rm{H}+}_{2}=\lambda^{\rm{H}-}_{2}=0$, and define the unipolarity parameters as $\lambda^{\rm{U}+}_{2}=1$, $\lambda^{\rm{U}-}_{2}=1/50$. This creates a unipolarity ratio $\lambda^{\rm{U}+}_{2}/\lambda^{\rm{U}-}_{2}=50$. By setting identical unipolarity and homogeneity parameters we hope to encourage QUACK to provide equal consideration to homogeneity in EUV and magnetic unipolarity. We note, however, that initial experiments performed over CR 2133 at a reduced spatial resolution suggest that if $\lambda^{\rm{H}+}_{1}=\lambda^{\rm{U}+}_{2}$ and $\lambda^{\rm{U}+}_{2}/\lambda^{\rm{U}-}_{2}\geq 10$, the specific choice of $\lambda^{\rm{U}-}_{2}$ has negligible effect.
@@
Figure 1: Comparison between full-scale segmentation results using only 193 Å data (ACWE), and using both 193 Å and HMI magnetogram data (QUACK), organized by CR. The box outlines the range between Q1 and Q3, with the median value in orange. The whiskers show 1.5 times the interquartile range. Outliers are marked with circles.
The IOU, SSIM, GCE, and LCE between the full-scale 193 Å ACWE segmentations of Grajeda et al. (2023) and QUACK, organized by CR, are presented in Figure 1. We note that, across all four metrics, CR 2133 displays the largest median discrepancy. Given that filaments are most prevalent in CR 2133, this is consistent with the expectation that the magnetic unipolarity parameters are reducing filament contamination. Visual inspection of the segmentation pairs with the lowest IOU for CR2133, of which three samples are provided in Figure 2, confirms this speculation by showing that filament regions, while still present, are consistently diminished to a small region surrounding the initial seed.
@@
Figure 2: Examples of full-scale segmentations using only 193 Å (ACWE, center), and using both 193 Å and HMI (QUACK, right) over CR 2133. The leftmost image is the 193 Å observation. The title of each figure is the record time for the EUV data.
Referring to Figure 1, we note that that CR 2099 shows a large range between Q1 and Q3 in IOU, GCE, and LCE, but not in SSIM. We also note that CR 2099 shows the second largest median discrepancy in IOU, but not in the remaining three metrics. These results are consistent with a change in overall CH area while retaining the overall form or shape of the identified CHs. Grajeda et al. (2023) note that CR 2099 had the largest number of segmentations that changed from targeting CHs to QS. This may suggest that QUACK is reducing the growth of $C^{+}$ into QS regions, which is another motivation for inclusion of magnetic field into evolution rather than just post-hoc processing. Visual inspection of the segmentation pairs with the lowest IOU for CR 2099, of which three samples are provided in Figure 3, also confirms this.
@@
Figure 3: Examples of full-scale segmentations using only 193 Å (ACWE, center), and using both 193 Å and HMI (QUACK, right) over CR 2099. The leftmost image is the 193 Å observation. The title of each figure is the record time for the EUV data.
In summary, incorporating magnetic field information, via a unipolarity metric operating on HMI data, appears to have negligible impact on segmentation of CHs. At the same time, it appears to minimize area of detected filaments and alleviate many cases of change of target. We return to these results in Section 4 after first considering effects of spatial resolution on segmentation accuracy.
@@ 3.4 Effects of Spatial Resolution on Segmentation Accuracy
The iterative nature of ACWE results in segmentations at full-scale taking several minutes, while reducing the image to one-eighth-scale resolution ($512\times 512$ pixels) reduces segmentation time to seconds (Grajeda et al., 2023). To determine the viability of using one-eighth-scale resolution images for QUACK segmentations, we replicate the pipeline of Grajeda et al. (2023) for reduced scale segmentations. We reduce the 193 Å image and corresponding magnetogram to $512\times 512$ pixels using scikit-image function skimage.transform.resize (van der Walt et al., 2014) after aligning the observations, but before correcting for limb brightening and before defining the seed. We evolve the contour on the decimated images, using the same weights as Section 3.3. After each segmentation has converged, it is scaled back to $4096\times 4096$ pixels and compared with the QUACK segmentations of Section 3.3.
@@
Figure 4: Comparison between full-scale and one-eighth-scale QUACK segmentations for CRs 2099-2010 (left) and CR 2133 (right). The box outlines the range between Q1 and Q3, with the median value in orange. The whiskers show 1.5 times the interquartile range. Outliers are marked with circles.
The effects of reducing spatial resolution on segmentation similarity are summarized in Figure 4. In this figure, we provide similarity for bi-linear interpolation, noting that five additional interpolation methods (nearest-neighbor, bi-quadratic, bi-cubic, bi-quartic, and bi-quintic) yielded similar results. Results for CRs 2099-2101 (left in each plot) are consistent with Grajeda et al. (2023), and results for CR 2133 (right in each plot) are improved. This suggests that decimation to one-eighth-scale is a viable means to reduce algorithm runtime despite (and in some cases, because of) the decimated magnetogram data.
Visual inspection revealed the same three discrepancies between full- and reduced-resolution segmentations as in Grajeda et al. (2023): 1) absence of smaller regions due to downsampling removing dark pixels that formed the initial seed, 2) presence or absence of spurious bright regions at different scales, and 3) reduced fidelity along the contour boundary due to the lower spatial resolution. All three effects are demonstrated in Figure 5. No new discrepancies appear to have been added by introducing spatially downsampled magnetogram data.
Figure 5: Example of full-scale and one-eighth-scale QUACK segmentations. The title of the figure is the record time of the EUV data.
Absence of smaller regions due to downsampling is especially prevalent in the case of filaments, which, due to their thin structure, are more likely to lose dark pixels that form the initial seed when downsampled. This explains the larger discrepancy Grajeda et al. (2023) found when rescaling CR 2133. In particular, the fact that adding magnetogram data constrains filament evolution at full-scale ensures higher similarity between filament-absent one-eighth-scale segmentations and filament-present full-scale segmentations.
@@
Figure 6: Comparison between full-scale and one-eighth-scale QUACK segmentations for CR 2133 when the full-scale seed is utilized to generate the one-eighth-scale segmentation. The box outlines the range between Q1 and Q3, with the median value in orange. The whiskers show 1.5 times the interquartile range. Outliers are marked with circles.
@@
Figure 7: Effects of seed transfer on QUACK segmentations generated at one-eighth-scale (right) compared to the corresponding full-scale segmentations (center). In each figure the leftmost image is the 193 Å observation and the title is the record time for the EUV data.
@@
Due to the aforementioned absence of some filaments in CR 2133 at one-eighth-scale, we additionally studied the effects of transferring the seed (initial contour) generated at full-scale resolution to the one-eighth scale resolution. This is done to ensure that all filaments present in full-scale segmentations are also in the one-eighth scale segmentations, and therefore to verify that spatially-decimated magnetogram data still effectively constrain CH segmentations. Due to differing spatial resolutions ($4096\times 4096$ versus $512\times 512$ pixels), the full-scale seed must be downsampled to the new resolution. Since every pixel at the one-eighth-scale resolution represents an $8\times 8$ pixel patch of the original image, we apply dilation with a $4\times 4$ pixel square using scikit-image function skimage.morphology.dilation (van der Walt et al., 2014) prior to downsampling the initial seed. This ensures that a region represented with even only one pixel in the full-scale seed is retained at the one-eighth-scale resolution. The effects of the seed transfer are summarized in Figure 6. We note that seed transfer had minimal effect on SSIM, GCE, and LCE, but further improved IOU, suggesting that the added filaments were constrained in the same manner as before. This is confirmed through a visual inspection of the data. For reference, Figure 7 provides full-scale and one-eighth-scale segmentations generated utilizing the seed transfer process for the same examples in Figure 2. Transferring the seed for the remaining three CRs had no appreciable effect.
@@
In summary, reducing the resolution of HMI magnetograms via standard image resizing procedures does not appear to affect segmentation performance. We note that this image decimation does not take into account magnetic flux conservation (i.e., correction for differing areas of pixels across the disk), but appears to retain statistics relevant to measures of unipolarity. Additionally, we find that a loss of seeds (individual regions in the initial contour) related to filaments does not appear to be the largest contributing source of this performance. Rather, it appears that reduced-resolution HMI data retains relevant information for constraining evolution of the contour, reducing the contribution of filaments. Since filaments are still present, albeit with reduced area, we now consider incorporation of magnetic field information into the seeding process to further reduce contamination. This reinforces the importance of incorporating magnetic field information throughout CH segmentation rather than as a post-hoc process.
@@ 4 Incorporation of Magnetic Field Data into the Seeding Process
Due to reduced filament contamination, improved runtime, and minimal effect on segmentation, we operate at one-eighth scale for all steps in QUACK. We recognize, however, that while this improves the quality of the segmentations, filament contamination, though minimized, persists. This section, therefore, explores the use of magnetic field data in initial seeding to further reduce contamination.
@@ 4.1 Complications with Estimating Unipolarity from Initial Seeds
Predicting the unipolarity of a final region through direct evaluation of the seed, regardless of the scale at which seeding occurred, is not possible due to two issues. First, multiple seeds (individual, spatially unconnected regions in the seed) often contribute to a single CH (these seeds merge into the final region during contour evolution). Second, these seeds are often small regions vulnerable to the effects of small-scale statistics, which results in erroneous estimations of unipolarity at the seeding stage. In order to demonstrate these two issues, we briefly return to ACWE (EUV-only) segmentations for the following two studies.
@@
First, regardless of scale, multiple regions within the initial seed often contribute to a single CH or filament. The effects of this are seen in Figure 2, where the use of magnetogram data minimized evolution of filament regions, resulting in small areas around the initial seeds rather than the unified regions seen in ACWE segmentations. Further demonstration of this behavior is shown by counting the number of regions in the seed and comparing that to the number of regions in the final ACWE segmentation. This comparison for CR 2133 is provided for full-scale (Figure 8) and one-eighth-scale (Figure 9) ACWE segmentations. In both figures the diagonal line represents the case where one region (CH or filament) is produced in the final segmentation for each region in the seed. Any case above this line indicates more regions in the final segmentation (region splitting), while any case below the line indicates more regions in the initial seed compared to the final contour (region merging). We note that region splitting does not occur, however region merging is prevalent.
@@
Figure 8: Scatter plot showing the number of regions in the initial seed (x-axis) and in the final segmentation (y-axis) for EUV-only (ACWE) segmentation at full scale for CR 2133. The color of each point denotes the number of images with that specific number of input and output regions. The diagonal line is the $x=y$ line, representing the case where there is one region in the final segmentation for every region in the initial seed.
@@
Figure 9: Scatter plot showing the number of regions in the initial seed (x-axis) and in the final segmentation (y-axis) for EUV-only (ACWE) segmentation at one-eighth scale for CR 2133. The color of each point denotes the number of images with that specific number of input and output regions. The diagonal line is the $x=y$ line, representing the case where there is one region in the final segmentation for every region in the initial seed.
@@
Multiple seeds contributing to a single CH is only problematic in that it exacerbates the second issue: small-scale statistics. It is reasonable to assume that regions in the initial contour with close proximity are part of the same final region, however, we note that regions with a large distance between them can belong to the same solar feature. In the counts for Figures 8 and 9, we account for proximity as we did in Grajeda et al. (2023). This is done by defining a region as all pixels that will become a single eight-connected component when dilated with a $40\times 40$ pixel square at full scale, or a $5\times 5$ pixel square at one-eighth scale. This same definition is applied in Figure 10 (full scale) and Figure 11 (one-eighth scale) which show scatter plots of initial versus final unipolarity for each region in the ACWE segmentations.
@@
Figure 10: Scatter plot showing the LOS unipolarity of seeds (x-axis) and corresponding regions in the final segmentation (y-axis) for EUV-only (ACWE) segmentation at full scale for all seed regions in CR 2133.
@@
Figure 11: Scatter plot showing the LOS unipolarity of seeds (x-axis) and corresponding regions in the final segmentation (y-axis) for EUV-only (ACWE) segmentation at one-eighth-scale for all seed regions in CR 2133.
@@
We note that both Figures 10 and 11 indicate regions with a high final (y-axis) value across all initial (x-axis) unipolarity values. This indicates that bipolar regions may appear to have any unipolarity, including high unipolarity, when evaluating the seed. We also note, referring to the cluster along the y-axis of Figure 10, that regions reporting high unipolarity in the seed are especially prevalent at the full-scale resolution. This explains why the initial seeds in filament regions remain, even when evolving using both EUV and magnetogram data in QUACK. The significantly smaller cluster on the y-axis of Figure 11 provides further indication that seeding at one-eighth-scale resolution helps minimize filament contamination. It should also be stated that the results in both Figures 10 and 11 contain all seeds across all 618 observations in CR 2133. For this reason, the same CH region is represented multiple times: once for every seed in every observation, as observed from one hour to the next. The spread, specifically along the x-axis of both figures, further indicates that subtle changes to the seed can strongly affect the apparent unipolarity of a region when evaluating the seed. These subtle changes to the seed are due to a handful of pixels being included or excluded due to minute variations in EUV intensity from one hour to the next and exacerbate the issue of inaccurate initial polarity.
@@
In summary, multiple seeds may contribute to a final evolved region (CH or filament). Furthermore, seeds may provide an inaccurate estimate of the unipolarity of the final region, mainly due to small-sample statistics. In light of these results, we consider a means to remove regions associated with filaments.